# Reconnaissance with JWST of the J-region Asymptotic Giant Branch in Distance Ladder Galaxies: From Irregular Luminosity Functions to Approximation of the Hubble Constant

### 09 Jan 2024 v1

@@ Abstract
We study stars in the J-regions of the asymptotic giant branch (JAGB) of near-infrared color magnitude diagrams in the maser host NGC 4258 and 4 hosts of 6 Type Ia supernovae (SN Ia): NGC 1448, NGC 1559, NGC 5584, and NGC 5643. These clumps of stars are readily apparent near and =22-25 mag with James Webb Space Telescope NIRCam photometry. Various methods have been proposed to assign an apparent reference magnitude for this recently proposed standard candle, including the mode, median, sigma-clipped mean or a modeled parameter of its luminosity function. We test the consistency of these by measuring intra-host variations, finding differences of up to 0.2 mag that significantly exceed statistical uncertainties. Brightness differences appear intrinsic, and are further amplified by the non-uniform shape of the JAGB LF, an issue already apparent in comparing the LMC and SMC. We follow a “many methods” approach to consistently measure JAGB magnitudes and distances to the SN Ia host sample calibrated by NGC 4258. We find broad agreement with the distances measured from Cepheids, tip of the red giant branch (TRGB) and Miras. However, the SN host mean distance estimated via the JAGB method and necessary to estimate , differs by 0.18 mag amongst the above definitions, a result of the different levels of asymmetry of the JAGB luminosity functions. The methods yield a full range of km s Mpc, i.e., a fiducial result of (stat) 2.3 (sys) ( 3.2 if combined in quadrature) km s Mpc, independent of Cepheids, TRGB, or Miras, with systematic errors limited by the differences in methods. Future work may seek to further standardize and refine this promising tool, making it more competitive with established distance indicators.
@@ 1 Introduction
The need for primary distance indicators that can reach tens of Mpc motivates efforts to identify alternative luminous standard candles, such as the recently proposed stars in the J-region Asymptotic Giant Branch (JAGB). The JAGB can be identified as an enhanced density of stars or a “clump” in the near-infrared color magnitude diagram (CMD; often in vs. or a similar filter set) redward of the red giant branch (RGB) and brighter than the tip of the RGB . This region is labelled ‘J’ in Fig. 1, corresponding to the labels defined by Weinberg & Nikolaev (2001), and is expected from stellar theory to be populated by thermally-pulsing carbon-rich (photospheric carbon-to-oxygen ratio ), dust-producing asymptotic giant branch stars with estimated amplitudes of 0.7 mag (Madore & Freedman, 2020) and J-band luminosities of 6 mag. Various measures of the apparent magnitude of the JAGB region, such as the mode, mean, median or a fitted value of its luminosity function, have been suggested and used in the literature as a potential standard candle that is luminous enough to measure extragalactic distances at megaparsec scales (Battinelli & Demers, 2005; Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Lee et al., 2021b; Zgirski et al., 2021; Parada et al., 2021; Madore et al., 2022; Parada et al., 2023; Lee, 2023).
@@
The theoretical foundation for expecting the mean of the JAGB to be a useful standard candle is rooted in the prediction that only a narrow range of stellar masses allows for a large quantity of carbon to be transported to the star’s surface, which results in the formation of carbonaceous circumstellar dust. These stars are expected to be of intermediate age and metallicity, with ages ranging from 300 Myr to 1 Gyr and masses ranging from 2 to 5 (Marigo et al., 2008; Karakas, 2014; Madore & Freedman, 2020), although these ranges may change with metallicity of the stars (Karakas, 2014). When a star’s mass falls within the prescribed range, the third dredge-up events during post-main sequence mass transport lead to the movement of carbon to the surface, resulting in the formation of a carbon AGB star (Groenewegen & Marigo, 2004). Stars with too little mass () do not have dredge up events that are efficient enough to bring enough carbon to the surface of the star to form a carbon star. Stars with too much mass will experience hot bottom burning, where carbon burns at the bottom of the convection zone before it can reach the surface. Observationally, the C/O1 feature of carbon stars is characteristically shown by carbon spectral bands such as CN, C, and CO, and lack of oxides (see for instance Gonneau et al., 2017, and references therein).
@@
Previous studies have indicated that the mass range required for an oxygen-rich asymptotic giant branch (AGB) star to transition into a carbon-rich AGB star can vary with the metallicity of the stellar population (Karakas, 2014; Ripoche et al., 2020; Parada et al., 2021, 2023). Lee (2023) showed that the JAGB does not show any dependency of its luminosity with metallicity over a very small explored range in [M/H] of 0.08 dex (). However, due to the small range in metallicity studied, we caution that their data is simultaneously consistent with no trend and with what could be a very large trend of 1 mag/dex. To show this, we show in Fig. 2 the data points from the reddening corrected JAGBs in Fig. 8 of Lee (2023); the nominal best-fit trend (red line) has a slope of mag/dex.111We also show the Spearman’s coefficient in the bottom left hand corner, which matches the value found in Lee (2023). The limited leverage due to the small range of metallicity results in a constraint that is consistent to within 1- with both no metallicity dependence and what would be a very large metallicity dependence of 1 mag/dex (cyan line in Fig. 2). For reference, we also plot a line with slope of 0.2 mag/dex, corresponding to approximately the most widely found Cepheid metallicity dependence (Breuval et al., 2022; Trentin et al., 2023, and references therein). All three of the lines shown in Fig. 2 are consistent with the data presented in Lee (2023), indicating no strong constraint. For context, if the metallicity dependence was as large as 1 mag/dex, the effort to calibrate to 5% precision without any metallicity correction would restrict the measurement to a field about 0.3 arcminutes wide at the distances of SN Ia host galaxies, based on typical host metallicity gradients of 0.05 dex/kpc (Hoffmann et al., 2016).
@@
Another issue which challenges the use of JAGB as a new standard candle derives from the change in shape (or population mix) indicated in its luminosity function between different hosts. Parada et al. (2021) show significantly different levels of asymmetry in the JAGB LFs in different galaxies: the LMC and NGC 6822 exhibit highly skewed luminosity functions, in contrast to the symmetric luminosity functions in the SMC and IC 1613; these differences are not due to foreground or background contamination. In the Local Group, the asymmetric LF in the LMC results produces a difference between its mode and mean of 0.15 mag, while this difference for the SMC is only mag. The magnitude estimates based on the median or a tightly-framed mean are in good agreement with their relative geometric distances; the distances based on the mode are not (Parada et al., 2021). We will revisit this point in the Discussion. Tight-framing of the mean (i.e., the selection of a small range to average), as employed by Madore & Freedman (2020) and Freedman & Madore (2020), can be used to avoid population contamination, and may offer good agreement with other distance indicators, but the lack of a defined prescription for these tight frames makes it difficult to evaluate and replicate this procedure. The non-trivial width of the J-region LF ( 0.3 mag for a single epoch and 0.2 mag using time-averaged magnitudes) also suggests some diversity of the stellar populations in this region. Despite these issues, the JAGB LF still offers one of the few possible standard candles luminous enough to rival the distance reach of Cepheids, TRGB, and Miras (Lee et al., 2023), motivating the present empirical study.
We find the JAGB clumps are readily apparent in the CMDs measured from JWST Cycle 1 program (GO-1685, P.I.: A. Riess, Riess et al., 2021), which observed four SN Ia hosts within 25 Mpc and NGC 4258 (obtained to measure Cepheids, Miras, and the TRGB). This further motivated us to investigate the feasibility of using the JAGB as a standard candle following the methods employed in the literature. We first investigate variations in the measured JAGB within hosts. We then measure distances and following a distance ladder and “many methods” approach. In Sections 2.1 to 2.3 we describe the data processing used for this study. We introduce how we measure the JAGB in Section 2.4. In Section 3 and 3.3, we investigate the JAGB through null tests, by varying the region of the measurement within a host. We characterize the variations in the JAGB measurements and then use them to determine distances to the SN Ia hosts in Section 4. We discuss our results in Section 5.
@@ 2.1 Data Selection
We provide a summary of the observations analyzed in this study in Table 1 and show their footprints in Fig. 3. For all JAGB measurements presented below, we correct for foreground extinction by retrieving values using the dustmaps package (Green, 2018) in Python and galaxy coordinates from the NASA/IPAC Extragalactic Distance Database (NED)333https://ned.ipac.caltech.edu/. The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. and applying the extinction coefficients listed in Table 3 assuming =3.1 and = 0.6021, and 0.2527 for F150W, and F277W, respectively. We measure all JAGBs in the main body of this paper in F150W.
@@ 2.2 DOLPHOT Photometry
We measure PSF photometry using the NIRCam version of DOLPHOT (Dolphin, 2000, 2016) in the Vega-Vega system. For this analysis, we used the beta version of the NIRCam module of DOLPHOT released on April 6, 2023. The most recent update as of writing is from December 2, 2023; this includes two changes since the April 6, 2023 version that we do not expect to significantly affect our results. The first change was the inclusion of the ‘Sirius-Vega’ zero-point. In this paper, we choose to remain with the original ‘Vega-Vega’ zero-point. The second change incorporates the “etctime” option, which allows the user to adjust the exposure times provided in the image headers. The choice of exposure times do not affect the photometry, although it may affect the photometric errors estimated with Poisson statistics. The photometric errors for our observations are in general too small to have a significant impact on the JAGB luminosity function, which has a typical width of 0.2 mag or more. See also Anand et al. (in prep.) for a discussion of these features. In addition, any future updates will likely minimally affect the results presented here because our measurements are all relative (i.e., between SN Ia hosts and NGC 4258 or within a host).
@@ 2.3 Crowding Corrections
where is the magnitude of the JAGB stars after crowding correction, is the observed magnitude of the same star, is the slope of the fit between the crowding bias and magnitude, and is the intercept of that same fit. We use the Rockfish cluster at Advanced Research Computing at Hopkins (ARCH) together with GNU parallel (Tange, 2011) to significantly reduce the time needed to run these artificial star measurements. In addition to determining the crowding bias, we also use the artificial star test results for NGC 5584, the furthest of the galaxies analyzed here, to test for possible incompleteness of the photometry. At the faintest test point of 1 mag fainter than 24.9 mag (25.9 mag) in F150W for NGC 5584 we recover approximately 97% of artificial stars in the region used to measure the JAGB. We provide the photometry used to measure the JAGB after crowding correction and crossmatching in the Github repository (upon publication), https://github.com/siyangliastro/JWST_JAGB_GO1685, and the CMDs using this photometry in Fig. 5.
@@ 2.4 Selection of a JAGB Reference Magnitude
In Fig. 4, we show CMDs for NGC 4258 containing the J-region in F150W vs. (see Fig. 3) on the West (chips A3+A4) and East (chips B3+B4) sides of the galaxy. We will refer to these two regions as the ‘Inner’ regions, as they appear to be on the inner disk and they are inside the “Outer” chips, A1, A2, B1, B2, at or near the halo. JAGB studies in the literature typically apply color and magnitude cuts to isolate the enhanced density of stars in the J-region (see, for instance, Zgirski et al. 2021); we approximate past efforts to put a “box” around the JAGB clump by defining a color cut of of 1 to 1.5 mag, and for NGC 4258 we use magnitude cuts of F150W = 21 to 24 mag (spanning about 5 standard deviations, i.e., a broad frame). This is the region marked as a blue dashed box in Fig. 4. We enlarge the points in the blue dashed box for emphasis. A binned luminosity function corresponding to stars inside the box can be seen in the right hand side of each subplot.
@@
To illustrate a specific prescription, we can use the mode-based method from Lee et al. (2023), which bins the luminosity function in 0.01 mag widths, applies Gaussian-windowed, locally weighted scatterplot smoothing (GLOESS; Loader, 2004; Persson et al., 2004) to the binned luminosity function with a smoothing scale of 0.25 mag, and uses the mode of the smoothed luminosity function to define the reference magnitude of the JAGB with an uncertainty set to the standard deviation of the LF divided by the square root of the number of stars in contains. We show the smoothed luminosity function for these measurements in blue and the corresponding mode-based JAGB measurement with the red line and label in Fig. 4.
@@
Unfortunately, the luminosity functions we observed for the JAGB in NGC 4258 are quite asymmetric (skew 0.4) which makes the selection of the measure of the reference magnitude non-trivial because its value will differ substantially depending on the choice of method, and this choice will not cancel when we compare to other LFs if their shapes are not the same. The asymmetry seen here is not a fluke or the result of incompleteness or extinction; in fact, a varying degree of one-sided asymmetry appears to be intrinsic to JAGB luminosity functions. As previously noted, asymmetric examples for nearby galaxies include the LMC and NGC 6822 (skew 0.3-0.5), while symmetric examples include the SMC and IC 1613 (skew , Parada et al., 2021).
@@
Depending on the level of LF asymmetry, we note that even the measurement of its mode will vary depending on the smoothing scale used as illustrated in the Appendix A, giving us an additional choice to consider. It is also far from obvious which measure of the LF will yield the better standard candle; this depends on whether the breadth of the LF contains multiple populations, the most common of which may be the standard candle, or whether it results from the spreading of a population whose mean is the true candle. Literature studies generally choose one method but none have provided a systematic study demonstrating the superiority of one or another.555Although as previously noted, the mean and median are preferred over the mode for comparing JAGB and the geometric distances for the LMC and SMC which we review in the Discussion. Therefore we will explore several different measures and retain a “many methods” approach to comparing these.
@@
We use the following four measures of the JAGB LF: median, 3-sigma clipped mean, mode, and GaussianQuadratic model fit as used (or modified) from recent JAGB literature (Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Zgirski et al., 2021; Parada et al., 2021; Lee et al., 2021b, 2022, a; Parada et al., 2023; Lee, 2023). We use 3 clipping to the mean to reduce sensitivity to outliers and box boundaries, features already inherent with the median and mode. In addition, we use the simple median instead of the modified median combined with skew from Ripoche et al. (2020); Parada et al. (2021, 2023). We will also vary the color range used to select the J-region stars in five different ways. In all we will use 25 combinations (5 measures, 5 color ranges) as listed in Table 6 to provide a comprehensive view.

### 29 Feb 2024 v2

@@ Abstract
We study stars in the J-regions of the asymptotic giant branch (JAGB) of near-infrared color magnitude diagrams in the maser host NGC 4258 and 4 hosts of 6 Type Ia supernovae (SN Ia): NGC 1448, NGC 1559, NGC 5584, and NGC 5643. These clumps of stars are readily apparent near and =22-25 mag with James Webb Space Telescope NIRCam photometry. Various methods have been proposed to assign an apparent reference magnitude for this recently proposed standard candle, including the mode, median, sigma-clipped mean or a modeled luminosity function parameter. We test the consistency of these by measuring intra-host variations, finding differences of up to 0.2 mag that significantly exceed statistical uncertainties. Brightness differences appear intrinsic, and are further amplified by the non-uniform shape of the JAGB luminosity function, also apparent in the LMC and SMC. We follow a “many methods’ approach to consistently measure JAGB magnitudes and distances to the SN Ia host sample calibrated by NGC 4258. We find broad agreement with distances measured from Cepheids, tip of the red giant branch (TRGB), and Miras. However, the SN host mean distance estimated via the JAGB method necessary to estimate differs by 0.19 mag amongst the above definitions, a result of different levels of luminosity function asymmetry. The methods yield a full range of km s Mpc, i.e., a fiducial result of (stat)2.3(sys)(3.1 if combined in quadrature) km s Mpc, with systematic errors limited by the differences in methods. Future work may seek to further standardize and refine this promising tool, making it more competitive with established distance indicators.
@@ 1 Introduction
The need for primary distance indicators that can reach tens of Mpc motivates efforts to identify alternative luminous standard candles, such as the recently proposed stars in the J-region Asymptotic Giant Branch (JAGB). The JAGB can be identified as an enhanced density of stars or a “clump” in the near-infrared color magnitude diagram (CMD; often in vs. or a similar filter set) redward of the red giant branch (RGB) and brighter than the tip of the RGB. This region is labelled ‘J’ in Fig. 1, corresponding to the labels defined by Weinberg & Nikolaev (2001), and is expected from stellar theory to be populated by thermally-pulsing carbon-rich (photospheric carbon-to-oxygen ratio ), dust-producing asymptotic giant branch stars with estimated amplitudes of 0.7 mag (Madore & Freedman, 2020) and J-band luminosities of 6 mag. Various measures of the apparent magnitude of the JAGB region, such as the mode, mean, median or a fitted value of its luminosity function, have been suggested and used in the literature as a potential standard candle that is luminous enough to measure extragalactic distances at megaparsec scales (Battinelli & Demers, 2005; Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Lee et al., 2021b, a; Zgirski et al., 2021; Parada et al., 2021; Madore et al., 2022; Parada et al., 2023; Lee, 2023; Lee et al., 2024).
@@
The theoretical foundation for expecting the mean of the JAGB to be a useful standard candle is rooted in the prediction that only a narrow range of stellar masses allows for a large quantity of carbon to be transported to the star’s surface, which results in the formation of carbonaceous circumstellar dust. These stars are expected to be of intermediate age and metallicity, with ages ranging from 300 Myr to 1 Gyr and masses ranging from 2 to 5 (Marigo et al., 2008; Karakas, 2014; Madore & Freedman, 2020), although these ranges may change with metallicity of the stars (Karakas, 2014). When a star’s mass falls within the prescribed range, the third dredge-up events during post-main sequence mass transport lead to the movement of carbon to the surface, resulting in the formation of a carbon AGB star (Groenewegen & Marigo, 2004). Stars with too little mass () (Groenewegen & Marigo, 2004) do not have dredge up events that are efficient enough to bring enough carbon to the surface of the star to form a carbon star. Stars with too much mass will experience hot bottom burning, where carbon burns at the bottom of the convection zone before it can reach the surface. Observationally, the C/O1 feature of carbon stars is characteristically shown by carbon spectral bands such as CN, C, and CO, and lack of oxides (see for instance Gonneau et al., 2017, and references therein).
@@
Previous studies have indicated that the mass range required for an oxygen-rich asymptotic giant branch (AGB) star to transition into a carbon-rich AGB star can vary with the metallicity of the stellar population (Karakas, 2014; Ripoche et al., 2020; Parada et al., 2021, 2023). Lee (2023) showed that the JAGB does not show any dependency of its luminosity with metallicity over a very small explored range in [M/H] of 0.08 dex (). However, due to the small range in metallicity studied, we caution that their data is simultaneously consistent with no trend and with what could be a very large trend of 1 mag/dex. To show this, we show in Fig. 2 the data points from the reddening corrected JAGBs in Fig. 8 of Lee (2023); the nominal best-fit trend (red line) has a slope of mag/dex.111We also show the Spearman’s coefficient in the bottom left hand corner, which matches the value found in Lee (2023). The limited leverage due to the small range of metallicity results in a constraint that is consistent to within 1 with both no metallicity dependence and what would be a very large metallicity dependence of 1 mag/dex (cyan line in Fig. 2). For reference, we also plot a line with slope of 0.2 mag/dex, corresponding to approximately the most widely found Cepheid metallicity dependence (Breuval et al., 2022; Trentin et al., 2023, and references therein). All three of the lines shown in Fig. 2 are consistent with the data presented in Lee (2023), indicating no strong constraint. For context, if the metallicity dependence was as large as 1 mag/dex, the effort to calibrate to 5% precision without any metallicity correction would restrict the measurement to a field about 0.3 arcminutes wide at the distances of SN Ia host galaxies, based on typical host metallicity gradients of 0.05 dex/kpc (Hoffmann et al., 2016).
@@
Another issue which challenges the use of JAGB as a new standard candle derives from the change in shape (or population mix) indicated in its luminosity function between different hosts. Parada et al. (2021) show significantly different levels of asymmetry in the JAGB LFs in different galaxies: the LMC and NGC 6822 exhibit highly skewed luminosity functions, in contrast to the symmetric luminosity functions in the SMC and IC 1613; these differences are not due to foreground or background contamination. In the Local Group, the asymmetric LF in the LMC results produces a difference between its mode and mean of 0.16 mag, while this difference for the SMC is only mag. The magnitude estimates based on the median or a tightly-framed mean are in good agreement with their relative geometric distances; the distances based on the mode are not (Parada et al., 2021). We will revisit this point in the Discussion. Tight-framing of the mean (i.e., the selection of a small range to average), as employed by Madore & Freedman (2020) and Freedman & Madore (2020), can be used to avoid population contamination, and may offer good agreement with other distance indicators, but the lack of a defined prescription for these tight frames makes it difficult to evaluate and replicate this procedure. The non-trivial width of the J-region LF ( 0.3 mag for a single epoch and 0.2 mag using time-averaged magnitudes) also suggests some diversity of the stellar populations in this region. Despite these issues, the JAGB LF still offers one of the few possible standard candles luminous enough to rival the distance reach of Cepheids, TRGB, and Miras (Lee et al., 2024), motivating the present empirical study.
We find the JAGB clumps are readily apparent in the CMDs measured from JWST Cycle 1 program (GO-1685, P.I.: A. Riess, Riess et al., 2021), which observed four SN Ia hosts within 25 Mpc and NGC 4258 (obtained to measure Cepheids, Miras, and the TRGB). This further motivated us to investigate the feasibility of using the JAGB as a standard candle following the methods employed in the literature. We first investigate variations in the measured JAGB within hosts. We then measure distances and following a distance ladder and “many methods” approach. In Sections 2.1 to 2.3 we describe the data processing used for this study. We introduce how we measure the JAGB in Section 2.4. In Section 3, we investigate the JAGB through null tests, by varying the region of the measurement within a host. We characterize the variations in the JAGB measurements and then use them to determine distances to the SN Ia hosts in Section 4. We discuss our results in Section 5.
@@ 2.1 Data Selection
We provide a summary of the observations analyzed in this study in Table 1 and show their footprints in Fig. 3. For all JAGB measurements presented below, we correct for foreground extinction by retrieving values using the dustmaps package (Green, 2018) in Python and galaxy coordinates from the NASA/IPAC Extragalactic Distance Database (NED)333https://ned.ipac.caltech.edu/. The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. and applying the extinction coefficients listed in Table 3 assuming =3.1 and = 0.6021 and 0.2527 for F150W, and F277W, respectively. We measure all JAGBs in the main body of this paper in F150W, noting that the filter is not a straight equivalent of the Two Micron All Sky Survey (2MASS) J-band that had been originally used for JAGB measurements such as in Madore & Freedman (2020). We discuss using the band for JAGB reference magnitude magnitudes and color relations in Appendix A.
@@ 2.2 DOLPHOT Photometry
We measure PSF photometry using the NIRCam version of DOLPHOT (Dolphin, 2000, 2016) in the Vega-Vega system. For this analysis, we used the beta version of the NIRCam module of DOLPHOT released on April 6, 2023. The most recent update as of writing is from December 2, 2023; this includes two changes since the April 6, 2023 version that we do not expect to significantly affect our results. The first change was the inclusion of the ‘Sirius-Vega’ zero-point. In this paper, we choose to remain with the original ‘Vega-Vega’ zero-point. The second change incorporates the “etctime” option, which allows the user to adjust the exposure times provided in the image headers. The choice of exposure times do not affect the photometry, although it may affect the photometric errors estimated with Poisson statistics. The photometric errors for our observations are in general too small to have a significant impact on the JAGB luminosity function, which has a typical width of 0.2 mag or more. See also Anand et al. (2024) for a discussion of these features. In addition, any future updates will likely minimally affect the results presented here because our measurements are all relative (i.e., between SN Ia hosts and NGC 4258 or within a host).
@@ 2.3 Crowding Corrections
where is the magnitude of the JAGB stars after crowding correction, is the observed magnitude of the same star, is the slope of the fit between the crowding bias and magnitude, and is the intercept of that same fit. We use the Rockfish cluster at Advanced Research Computing at Hopkins (ARCH) together with GNU parallel (Tange, 2011) to significantly reduce the time needed to run these artificial star measurements. In addition to determining the crowding bias, we also use the artificial star test results for NGC 5584, the furthest of the galaxies analyzed here, to test for possible incompleteness of the photometry. At the faintest test point of 1 mag fainter than 24.9 mag (25.9 mag) in F150W for NGC 5584 we recover approximately 97% of artificial stars in the region used to measure the JAGB. We provide the photometry used to measure the JAGB after crowding correction and crossmatching in the Github repository (upon publication), https://github.com/siyangliastro/JWST_JAGB_GO1685, and Zenodo repository, 10.5281/zenodo.10666761 (catalog DOI) 10.5281/zenodo.10666761, and the CMDs using this photometry in Fig. 5.
@@ 2.4 Selection of a JAGB Reference Magnitude
In the top row of Fig. 4, we show CMDs for NGC 4258 containing the J-region in F150W vs. (see Fig. 3) on the West (chips A3+A4) and East (chips B3+B4) sides of the galaxy. We will refer to these two regions as the ‘Inner’ regions, as they appear to be on the inner disk and they are inside the “Outer” chips, A1, A2, B1, B2, at or near the halo. JAGB studies in the literature typically apply color and magnitude cuts to isolate the enhanced density of stars in the J-region (see, for instance, Zgirski et al. 2021); we approximate past efforts to put a “box” around the JAGB clump by defining a color cut of of 1 to 1.5 mag, and for NGC 4258 we use magnitude cuts of F150W = 21 to 24 mag (spanning about 5 standard deviations, i.e., a broad frame). This is the region marked as a blue dashed box in Fig. 4. We enlarge the points in the blue dashed box for emphasis. A binned luminosity function corresponding to stars inside the box can be seen in the right hand side of each subplot.
@@
To illustrate a specific prescription, we can use the mode-based method from Lee et al. (2024), which bins the luminosity function in 0.01 mag widths, applies Gaussian-windowed, locally weighted scatterplot smoothing (GLOESS; Loader, 2004; Persson et al., 2004) to the binned luminosity function with a smoothing scale of 0.25 mag, and uses the mode of the smoothed luminosity function to define the reference magnitude of the JAGB. The uncertainty is set to the standard deviation of the LF divided by the square root of the number of stars in contains. We show the smoothed luminosity function for these measurements in blue and the corresponding mode-based JAGB measurement with the red line and label in Fig. 4.
@@
Unfortunately, the luminosity functions we observed for the JAGB in NGC 4258 are quite asymmetric (skew 0.4 for the Inner fields) which makes the selection of the measure of the reference magnitude non-trivial because its value will differ substantially depending on the choice of method, and this choice will not cancel when we compare to other LFs if their shapes are not the same. The asymmetry seen here is not a fluke or the result of incompleteness or extinction; in fact, a varying degree of one-sided asymmetry appears to be intrinsic to JAGB luminosity functions. As previously noted, asymmetric examples for nearby galaxies include the LMC and NGC 6822 (skew 0.4 to 0.5), while symmetric examples include the SMC and IC 1613 (skew , Parada et al., 2021).
@@
Depending on the level of LF asymmetry, we note that even the measurement of its mode will vary depending on the smoothing scale used as illustrated in the Appendix B, giving us an additional choice to consider. It is also far from obvious which measure of the LF will yield the better standard candle; this depends on whether the breadth of the LF contains multiple populations, the most common of which may be the standard candle, or whether it results from the spreading of a population whose mean is the true candle. Literature studies generally choose one method but none have provided a systematic study demonstrating the superiority of one or another555Although as previously noted, the mean is preferred over the mode for comparing JAGB and the geometric distances for the LMC and SMC which we review in the Discussion. Therefore we will explore several different measures and retain a “many methods” approach to comparing these.
@@
We use the following four measure types of the JAGB LF: median, 3-sigma clipped mean, mode, and GaussianQuadratic model fit as used (or modified) from recent JAGB literature (Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Zgirski et al., 2021; Parada et al., 2021; Lee et al., 2021b, 2022, a; Parada et al., 2023; Lee, 2023). We use 3 clipping to the mean to reduce sensitivity to outliers and box boundaries, features already inherent with the median and mode. In addition, we use the simple median instead of the modified median combined with skew from Ripoche et al. (2020); Parada et al. (2021, 2023). We will also vary the color range used to select the J-region stars in five different ways. In all we will use 25 combinations (5 measures, 5 color ranges) as listed in Table 6 to provide a comprehensive view.